Part I. Differentiation Rules on Algebraic Functions

Apply the appropriate differentiation technique/s to determine the derivative of the following algebraic functions.


  1. f(x)=x^(-6)+4x^3-10
  2. f(x)=5x^3-4/x^2
  3. y=√(5-2x)
  4. g(x)=(x^2-5x-7)(x^2-5)
  5. g(x)=(3x^2-4x+1)(2x-3)
  6. h(x)=(4x^3-2x+1)/(2x-3)
  7. h(x)=(4x+3)/(5x^2-3x+6 )
  8. y=(3x^2+2x-3)^3
  9. y=(5x+3)^2 (2x-1)
  10. y=(3x-2)^2/(4x+1)


Part II. Differentiation Rules on Trigonometric Functions and Higher-Order Derivatives

Apply the appropriate differentiation technique/s to determine the derivative of the following trigonometric functions.


  1. y=cos〗⁡(x^2+3)
  2. y=sin⁡〖(5x^2-7)〗
  3. y=cot⁡(6x)
  4. y=tan⁡(cos⁡x )
  5. y=csc⁡(5x^3 )


Find the first derivative, second derivative, and third derivative for each of the following functions.

  1. y=6x^4+3x^3-4x^2-1
  2. f(x)=(2x+3)^4
1
Expert's answer
2022-06-08T17:08:01-0400

Part I.

1.

"f'(x)=-6x^{-7}+12x^2"

2.

"f'(x)=15x^2+8\/x^3"

3.

"y'=-\dfrac{1}{\sqrt{5-2x}}"

4.

"g'(x)=(2x-5)(x^2-5)+(x^2-5x-7)(2x)"


"=2x^3-10x-5x^2+25+2x^3-10x^2-14x"

"=4x^3-15x^2-24x+25"

5.

"g'(x)=(6x-4)(2x-3)+2(3x^2-4x+1)"


"=12x^2-18x-8x+12+6x^2-8x+2"

"=18x^2-34x+14"

6.

"h'(x)=\dfrac{(12x^2-2)(2x-3)-2(4x^3-2x+1)}{(2x-3)^2}"


"=\dfrac{24x^3-36x^2-4x+6-8x^3+4x-2}{(2x-3)^2}"

"=\dfrac{16x^3-36x^2+4}{(2x-3)^2}"

7.

"h'(x)=\dfrac{4(5x^2-3x+6)-(10x-3)(4x+3)}{(5x^2-3x+6)^2}"


"=\dfrac{20x^2-12x+24-40x^2-30x+12x+9}{(5x^2-3x+6)^2}"

"=\dfrac{-20x^2-30x+33}{(5x^2-3x+6)^2}"

8.


"y'=3(3x^2+2x-3)^2(6x+2)"

"=6(3x+1)(3x^2+2x-3)^2"

9.


"y'=2(5x+3)(5)(2x-1)+2(5x+3)^2"

"=2(5x+3)(10x-5+5x+3)"

"=2(5x+3)(15x-2)"

10.

"y'=\dfrac{6(3x-2)(4x+1)-4(3x-2)^2}{(4x+1)^2}"


"=\dfrac{2(3x-2)(12x+3-6x+4)}{(4x+1)^2}"

"=\dfrac{2(3x-2)(6x+7)}{(4x+1)^2}"

Part II.

1.

"y'=-2x\sin(x^2+3)"

2.


"y'=10x\cos(5x^2-7)"

3.


"y'=-\dfrac{6}{\sin^2(6x)}=-6\csc^2(6x)"

4.


"y'=-\dfrac{\sin x}{\cos^2(\cos x)}=-\sin x\sec^2 (\cos x)"

5.


"y'=-\dfrac{1}{\sin^2(5x^3)}(\cos (5x^3))(15x^2)"

"=-15x^2\cot(5x^3)\csc x(5x^3)"

Part III.

1.


"y'=24x^3+9x^2-8x"

"y''=72x^2+18x-8"

"y'''=144x+18"

2.


"f'(x)=8(2x+3)^3"

"f''(x)=48(2x+3)^2"

"f'''(x)=192(2x+3)"